OpenStudy (anonymous):

If possible, choose k so that the following function is continuous on any interval (equation in post).

OpenStudy (anonymous):

\[f(x)=\left[\begin{matrix}kcos(x), & x \le0 \\ 10e^{x}-k,&0< x \end{matrix}\right]\]

OpenStudy (anonymous):

@rebeccaskell94

OpenStudy (anonymous):

I haven't passed 4th grade math yet.

OpenStudy (anonymous):

Oh that's ok then (= thanks anyhow

hartnn (hartnn):

do you know the basics of continuity ?

OpenStudy (anonymous):

that the line must contiously go and still be a function (one to one) =P if that's what you mean

hartnn (hartnn):

i meant, if f(x) is continuous, at x=0, \[\lim \limits{x \rightarrow 0^-} f(x)=\lim \limits{x \rightarrow 0^+f(x)}\]

hartnn (hartnn):

do you you something like that ^ ?

OpenStudy (anonymous):